OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation 4x - 3y = 24. My answers is x intercept = (8, 0); y intercept = (0, 6) Am I right? :)

OpenStudy (whpalmer4):

Here's how you can check: 4(8)-3(0) = 24 32 - 0 = 24 oops Here's how you identify the real x-intercept and y-intercept: x intercept is where y = 0. 4x-3(0) = 24 4x - 0 = 24 4x = 24 x = ? y intercept is where x = 0 4(0)-3y = 24 0-3y = 24 -3y = 24 y = ?